# InTAR: Inter-Task Auto-Reconfigurable Accelerator Design for High Data Volume Variation in DNNs

### v1

@@ Abstract
The rise of deep neural networks (DNNs) has driven a boom in AI services, which results in an increased demand for computing power and memory. In modern DNNs, the data sizes produced and consumed are highly varied across operations (high data volume variation, HDV). Because existing design paradigms use fixed execution patterns that lead to either low computational efficiency due to pipeline stalls or frequent off-chip memory accesses to manage large intermediate data, HDV applications are challenging to accelerate on FPGAs. To address these challenges, we introduce the Inter-Task Auto-Reconfigurable Accelerator (InTAR), a novel accelerator design for HDV applications on FPGAs. InTAR combines the high computational efficiency of sequential execution with the reduced off-chip memory overhead of dataflow execution. It switches execution patterns automatically with a static schedule determined before circuit design based on resource constraints and model parameters. Unlike previous reconfigurable accelerators, InTAR encodes reconfiguration schedules during circuit design, allowing model-specific optimizations that allocate only the necessary logic and interconnects. Thus, InTAR achieves a high clock frequency with fewer resources and low reconfiguration time. Furthermore, InTAR supports high-level tools such as HLS for fast design generation. We implement a set of multi-task kernels in various HDV DNNs using InTAR. Compared with dataflow and sequential accelerators, InTAR exhibits 1.8×1.8$\times$1.8 × and 7.1×7.1$\times$7.1 × speedups correspondingly. We also implement InTAR for GPT-2 medium as a more complex example, which achieves a speedup of $\mathbf{3.65\sim 39.14\times}$ and a $\mathbf{1.72\sim 10.44\times}$ boost in DSP efficiency compared to the corresponding SoTA accelerators (Allo and DFX) on FPGAs. Additionally, this design demonstrates $\mathbf{1.66\sim 7.17\times}$ better power efficiency than GPUs.
@@ I Introduction
The development of deep neural networks (DNNs) has driven numerous breakthroughs in AI-assisted applications [13, 14] while intensifying demands for compute efficiency due to increasing network complexity and memory requirements. As DNN applications grow in size [34, 30], larger data volumes generated during operations strain hardware resources during deployment. Consequently, a key challenge in designing DNN accelerators is: Given resource constraints imposed by physical hardware limitations or power considerations, how can execution scheduling be improved to minimize end-to-end latency while accounting for hardware overhead?
@@
  • Sequential: high off-chip memory access overhead. Output data from previous tasks will be stored and reused by the next tasks in sequential execution. Moreover, data may be cached and remain unused until the consumer tasks start executions. Due to the limited on-chip resources, it is often difficult to store output data on SRAM. If storing data in the off-chip memory, the accelerator will suffer from high latency and energy consumption of off-chip memory access.
@@
Therefore, we propose a novel accelerator design paradigm on FPGAs: inter-task auto-reconfigurable accelerator (InTAR). InTAR can switch execution patterns automatically based on on-chip memory and computation resources. When a task produces large intermediate data, InTAR pipelines multiple tasks to avoid accessing off-chip memory for these data. Otherwise, InTAR will process tasks sequentially to maximize compute efficiency by eliminating pipeline stalls. Compared with other reconfigurable accelerators [7, 10, 4, 26, 25], InTAR allows model-specific circuit optimization that keeps only necessary control logics and interconnects. Hence, InTAR requires fewer reconfiguration resources, achieves a high clock frequency, and low reconfiguration overhead (10 to 20 ns) (Section III). Since computations are reconfigured at the task level, InTAR is one of the first works regarding FPGA-based reconfigurable accelerators that support high-level hardware generation tools such as High-Level Synthesis (HLS) [11] for fast accelerator design. Our contributions include:
@@ II-A Dataflow and Sequential DNN Accelerators
Dataflow accelerators allocate resources for each task in the DNN and pipeline the computations to save memory while maintaining performance. FlexCNN [6] is a framework for composing CNN-based DNN accelerators in HLS, and GenGNN [3] is a framework for GNN acceleration. Allo [9, 8] is a composable framework for designing accelerators on FPGA by parsing Python/Pytorch scripts and generating the corresponding HLS code for FPGAs for every operator. Accompanied by proper customization primitives, Allo can automatically combine the optimizations across different layers for a complex model.
On the other hand, sequential accelerators process layers in DNN sequentially and attempt to utilize all available resources. GPUs and TPUs [24] are considered instruction-based sequential accelerators that support various DNNs. ZyncNet [18] is a CNN accelerator deployed on Zync SoC with a specialized topology. FQ-BERT [27] is a sequential FPGA design for a fully quantized BERT model [12]. FlightLLM [42] is a vector processor implemented on Alveo U280 FPGA for Llama model inference.
@@ II-B Hybrid Accelerators
Previous works attempt to mitigate the drawbacks of dataflow and sequential accelerators by having a hybrid execution pattern. A naive approach is allocating part of the resource for dataflow execution and the rest for sequential execution [43]. A more recent method is to allocate several PEs and schedule multiple tasks to the same PE. For instance, SSR [45] is a hybrid accelerator implemented on AMD Xilinx Versal ACAP devices for vision transformers. By reserving two groups of AI engines for two types of matrix multiplies, SSR allows sequential executions of GEMMs to minimize latency and enables data forwarding between dependent GEMMs to reduce the memory cost. However, PE specializations of hybrid accelerators impede the flexibility of execution pattern switching, which may lead to stalls. Section II-E will further discuss the inefficiency of hybrid accelerators with an example.
Figure 1: Example of mapping computations of attention and linear projection of value matrix to the dataflow, sequential, hybrid accelerators, and InTAR. Depending on the scheduling of the rest of the computations in the entire model, InTAR can choose between task-parallel (left) and sequential modes (right) for better locality. Both will have the same latency.
@@ II-C Reconfigurable Accelerators
A reconfigurable accelerator can change its microarchitecture on the fly. A common implementation of such accelerators is the coarse-grain reconfigurable array (CGRA) [26, 10, 31], which compiles a dataflow graph into a configuration and modifies switches connecting PEs accordingly. However, CGRA has complex reconfiguration logic to ensure generalizability. This introduces area overheads [32] and additional wiring that negatively affects the timing [26].
An alternative method to reconfigure resources is dynamic partial reconfiguration (DPR) [38], which reserves regions on FPGAs to load partial bitstream for runtime reconfiguration. While having a low resource overhead for the configuration controller, it has a higher reconfiguration overhead (10 to 100 ms) than FPGA-based CGRA (around 10 ns).
@@
Considering these drawbacks, existing reconfigurable accelerators can hardly apply to a broad range of modern DNNs, which necessitates the development of a new design paradigm.
@@ II-D Applications with High Data Volume Variation
A multi-task application is defined as high data-volume variation (HDV) if the input/output data sizes between tasks are highly varied. Many of DNNs are HDV for two reasons:
  • Modern DNNs have complex dependencies. Data is produced several layers ahead and can only be discarded after they are reused. Hence, the data production and elimination rates are largely varied.
  • The objectives of many DNN workloads involve feature extraction (e.g., classification) that reduces the data size, or data construction (e.g., image/text generation) which increases the data size. By the nature of these workloads, the data size will change during execution.
II-E Motivating Case Study: Attention + Linear Projection
We provide a case study to illustrate the advantages of inter-task reconfiguration for HDV DNNs: calculating the attention score and a linear projection of the value matrix. Figure 1 shows the dataflow graph of the attention layer and the timeline for each execution pattern. The direction of the arrows indicates the output production order and lengths represent the production rate. Here we focus on calculating $Q$, $K$, $V$, and $A$ in the dashed square. In this example, we assume that:
  • Matrix multiplies are weight stationary as weights are already stored on on-chip memory.
@@ II-E Motivating Case Study: Attention + Linear Projection
The pattern of data size variation is first increase then decrease. Dataflow execution instantiates PEs for each task. There are three ways to stream and cache data across tasks illustrated in Figure 1: outer product with aligned and unaligned production rates of $Q$ and $K$, and inner product with unaligned production rates. Inner products with aligned production rates require caching of both $Q$ and $K$, which is infeasible. Each case incurs compute resource idling due to either varied workload size of $A$ (Figure 1(a)), back-pressure (Figure 1(b)), or early completion (Figure 1(c)).
@@ III Inter-Task Auto-Reconfigurable Accelerator
Unlike other reconfigurable [26, 10, 33, 7, 25] and overlay accelerators [40, 41], which prioritize abstraction and fast compilation, InTAR focuses on reconfiguration for optimization. The reconfiguration schedule in InTAR is determined at circuit design time rather than relying on post-design external commands. For every application, a static schedule is created based on the specific application and resource constraints, and the circuit is then tailored accordingly. This approach reduces reliance on general-purpose control units and cross-bars, retaining only essential reconfiguration logic and interconnects. The result is low resource utilization for reconfiguration controls (49% LUT reduction for a design with 3×3$\times$3 × throughput compared to [26] for multiple GEMMs), high clock frequency (2.41×2.41$\times$2.41 × that of [26]), and low reconfiguration latency ($10^{6}\times$ faster than DPR [25], comparable to [26]) compared with FPGA-based CGRA and DPR approaches. Additionally, InTAR handles reconfiguration at the task level, enabling rapid and convenient development with high-level tools like HLS. This makes InTAR well-suited for FPGAs to adapt quickly to new DNN applications. Lastly, InTAR’s reconfiguration scope is not limited to specific computations. Table II compares InTAR with prior DNN accelerator works.
@@ III-A Execution Modes of InTAR
Figure 2: Left: Architecture template of InTAR. Compute cores (CC) compute linear operations (e.g., GEMM, ConvNet), and SFUs compute non-linear operations (e.g., softmax, GeLU). Each CC contains a scratchpad memory, a reconfigurable MAC unit array, a reduction unit, and a data movement control unit. Dashed lines indicate the candidates for connection between CCs and SFUs. Right: example architectures for each execution mode within the template, with $n,m=2$. For a schedule, we will compose the architectures of required modes to keep only necessary interconnects and logic.
@@
  • Task-pipeline: Same as dataflow execution with pipelined dependent tasks. Computational resources are partitioned proportional to the workloads. Data are streamed from one task to another using FIFOs.
  • Task-parallel: Same as dataflow execution with independent tasks running in parallel. The resource efficiency is the same as the sequential mode, but it may have a better locality for subsequent tasks (e.g., in Figure 1, if calculating $AV$ and $O$ are pipelined, then running $Q$ and $V$ in parallel and assigning $V$ to the PE that will compute $AV$ avoids data aggregation overhead from other PEs).
Figure 1(f) illustrates the task assignment and timeline of InTAR for the case study in Section II-E. PEs 1 and 2 handle both attention computation and linear projections. Depending on the subsequent operations, InTAR can choose either task-parallel or sequential mode for $Q$ and $V$. If the scheduler decides to serialize computing $AV$ and $O$, then the sequential mode is more suitable to allow PEs 1 and 2 to execute cooperatively. If the scheduler pipelines computations of $AV$ and $O$, then task-parallel mode is preferred to improve locality. Both choices will not affect the latency but determine the data movement overhead of the following tasks. After calculating $Q$ and $V$, InTAR switches to task-pipeline mode to compute $K$ and $A$, since storing $Q$ consumes all scratchpad memory.
@@ III-B Architecture Template
  • Reconfigurable PE Array: a multi-dimensional array of PEs that allow changing inputs and precisions. Connected in either a systolic array or broadcasting the data to all PEs. Each PE produces the complete or partial result of the linear operation.
@@
When switching execution modes, a global instruction reader will read both the configuration instructions from a static buffer and the input size passed as the kernel argument. It will then modify part of each instruction related to the input size (e.g., the loop bounds) and send it to the CCs. The reconfigurable PEA and data movement control will read the instructions and propagate among the CCs. The instruction sequence is determined by the static schedule of mode switching at the circuit design time. Each instruction contains a stage index and several loop bounds (e.g., reconfiguring the three loop bounds for GEMMs). A stage refers to an execution of a single or a group of tasks that follows one of the execution modes, seen in Figure 1(f), e.g., computing $Q$ and $V$ in stage 1 and $K$ and $A$ in stage 2. Each CC has a specified behavior for each stage, orchestrated by multiplexers with stage indices as the select signals. Notice that the stage-to-behavior correspondences are specific for each application. Thus, InTAR only needs to handle the reconfigurations defined in the schedule. The reconfiguration overhead consists of four parts: reading, modifying, sending, and decoding instructions to generate signals for multiplexers in the CCs. Each of these is a single-cycle operation. Thus, reconfigurations have a 4-cycle overhead (10 to 20 ns depending on frequency).
@@ IV-A Finding Schedule and Architecture Template Parameters
The development of InTAR for a new application starts with creating the static reconfiguration schedule. In this work, we first topologically sort the tasks based on dependencies in the dataflow graph and create the schedule based on the data size and resource constraints, using the following heuristics:
@@
Then, we analyze the generated schedule and the hardware configurations and determine architecture template parameters with the following heuristics: The schedule infers the instantiation of the SFUs and FIFO connections. The column count of CCs equals the maximum number of tasks of all stages, and the row count is the number of dies on the FPGA necessary to improve floorplanning. We compute the PEA dimensions to maximize utilization with limited cross-die communications and calculate the memory ports needed. Finally, the scratchpad memory is the maximum memory size required over all stages. Although the heuristics are not guaranteed optimal, we show that our schedule results improved performance consistently (Sections V, VI). We leave the development of an optimized DSE engine as future work.

### v2

@@ Abstract
The rise of deep neural networks (DNNs) has driven an increased demand for computing power and memory. Modern DNNs exhibit high data volume variation (HDV) across tasks, which poses challenges for FPGA acceleration: conventional accelerators rely on fixed execution patterns (dataflow or sequential) that can lead to pipeline stalls or necessitate frequent off-chip memory accesses. To address these challenges, we introduce the Inter-Task Auto-Reconfigurable Accelerator (InTAR)  [^1] https://github.com/OswaldHe/InTAR, a novel accelerator design methodology for HDV applications on FPGAs. InTAR combines the high computational efficiency of sequential execution with the reduced off-chip memory overhead of dataflow execution. It switches execution patterns automatically with a static schedule determined before circuit design based on resource constraints and problem sizes. Unlike previous reconfigurable accelerators, InTAR encodes reconfiguration schedules during circuit design, allowing model-specific optimizations that allocate only the necessary logic and interconnects. Thus, InTAR achieves a high clock frequency with fewer resources and low reconfiguration time. Furthermore, InTAR supports high-level tools such as HLS for fast design generation. We implement a set of multi-task HDV DNN kernels using InTAR. Compared with dataflow and sequential accelerators, InTAR exhibits $\mathbf{1.8\times}$ and $\mathbf{7.1\times}$ speedups correspondingly. Moreover, we extend InTAR to GPT-2 medium as a more complex example, which is $\mathbf{3.65\sim 39.14\times}$ faster and a $\mathbf{1.72\sim 10.44\times}$ more DSP efficient than SoTA accelerators (Allo and DFX) on FPGAs. Additionally, this design demonstrates $\mathbf{1.66\sim 7.17\times}$ better power efficiency than GPUs.
@@ I Introduction
The development of deep neural networks (DNNs) has driven numerous breakthroughs in AI-assisted applications [1, 2] while intensifying demands for compute efficiency due to increasing network complexity and memory requirements. As DNN applications grow in size [3, 4], larger data volumes generated between operations strain hardware resources during deployment. Consequently, a key challenge in designing DNN accelerators is: Given resource constraints imposed by physical hardware limitations or power considerations, how can execution scheduling be improved to minimize end-to-end latency while accounting for hardware overhead?
@@
  • Sequential: high off-chip memory access overhead. Output data from previous tasks will be stored and reused by the next tasks in sequential execution. Moreover, data may be cached and remain unused until the consumer tasks start executions. Due to the limited on-chip resources, it is often difficult to store all output data in SRAM. If storing data in the off-chip memory, the accelerator will suffer from high latency and energy consumption of off-chip memory access.
@@
Therefore, we propose a novel accelerator design paradigm on FPGAs: inter-task auto-reconfigurable accelerator (InTAR). InTAR can switch execution patterns automatically based on on-chip memory and computation resources. When a task produces large intermediate data, InTAR pipelines multiple tasks to avoid accessing off-chip memory for this data. Otherwise, InTAR will process tasks sequentially to maximize computational efficiency by eliminating pipeline stalls. Compared with other reconfigurable accelerators [17, 18, 19, 20, 21], InTAR allows model-specific circuit optimization that keeps only necessary control logics and interconnects for reconfiguration. Hence, InTAR requires fewer reconfiguration resources, achieves a high clock frequency, and has a low reconfiguration overhead (10 to 20 ns) (Section III). Moreover, since computations are reconfigured at the task level, InTAR is one of the first works regarding FPGA-based reconfigurable accelerators that support high-level hardware generation tools such as High-Level Synthesis (HLS) [22] for fast accelerator design. Specifically, our contributions include:
@@ II-A Dataflow and Sequential DNN Accelerators
Dataflow accelerators allocate resources per DNN task and pipeline computations to conserve memory while preserving performance. For example, FlexCNN [7] composes CNN-based accelerators in HLS, GenGNN [24] accelerates GNNs, and Allo [6, 12] designs FPGA accelerators by parsing Python scripts into HLS code for each operator, automatically combining layer optimizations with customization primitives.
On the other hand, sequential accelerators process layers in DNN sequentially and attempt to utilize all available resources. GPUs and TPUs [8] are considered instruction-based sequential accelerators that support various DNNs. ZyncNet [25] is a CNN accelerator deployed on Zync SoC with a specialized topology. FlightLLM [26] is a vector processor implemented on Alveo U280 FPGA for Llama model inference.
@@ II-B Hybrid Accelerators
Some previous works attempt to mitigate the drawbacks of dataflow and sequential accelerators by having a hybrid execution pattern. A naive approach is allocating part of the resource for dataflow execution and the rest for sequential execution [11]. A more recent method is to allocate several PEs and schedule multiple tasks to the same PE. For instance, SSR [13] is a hybrid accelerator implemented on AMD Xilinx Versal ACAP devices for vision transformers. By reserving two groups of AI engines for two types of matrix multiplies, SSR allows sequential executions of matrix multiplies (GEMM) to minimize latency and enables data forwarding between dependent GEMMs to reduce the memory cost. However, PE specializations of hybrid accelerators impede the flexibility of execution pattern switching, which may lead to stalls. Section II-E will further discuss the inefficiency of hybrid accelerators with an example.
Figure 1: Example of mapping computations of attention and linear projection of value matrix to the dataflow, sequential, hybrid accelerators, and InTAR. The sequence length is 256, the hidden dimension is 1024, and the weights are $1024\times 1024$ matrices. Depending on the scheduling of the rest of the computations in the entire model, InTAR can choose between task-parallel (left) and sequential modes (right) for better locality. Both will have the same latency.
@@ II-C Reconfigurable Accelerators
A reconfigurable accelerator can change its microarchitecture on the fly. A common implementation of such accelerators is the coarse-grain reconfigurable array (CGRA) [20, 18, 27], which compiles a dataflow graph into a configuration and modifies switches connecting PEs accordingly. However, CGRA has complex reconfiguration logic to ensure generalizability. This introduces substantial area overheads [28] and additional wiring that negatively affects the timing [20].
An alternative method to reconfigure resources is dynamic partial reconfiguration (DPR) [29], which reserves regions on FPGAs to load partial bitstreams for runtime reconfiguration. While having a low resource overhead for the configuration controller, it has a higher reconfiguration overhead (10 to 100 ms) than FPGA-based CGRA (around 10 ns).
@@
Considering these limitations, existing reconfigurable accelerators can hardly apply to a broad range of modern DNNs, which necessitates the development of a new design paradigm.
@@ II-D Applications with High Data Volume Variation
A multi-task application is defined as high data-volume variation (HDV) if the input/output data sizes between tasks are highly varied. Many of the DNNs are HDV for two reasons:
  • Modern DNNs have complex dependencies. Data is produced several layers in advance and can only be discarded after they are reused. Hence, the data production and elimination rates vary widely.
  • The objectives of many DNN workloads involve feature extraction (e.g., classification) that reduces the data size, or data construction (e.g., image/text generation), which increases the data size. By the nature of these workloads, the data size will change during execution.
II-E A Motivating Case Study: Attention + Linear Projection
We provide a case study to illustrate the advantages of inter-task reconfiguration for HDV DNNs: calculating the attention score and a linear projection of the value matrix. Figure 1 shows the dataflow graph of the attention layer and the timeline for each execution pattern. The direction of the arrows indicates the output production order, and the lengths represent the production rate. Here we focus on calculating $Q$, $K$, $V$, and $A$ in the dashed square. In this example, we assume that:
  • Matrix multiplies are weight-stationary as weights are already stored in on-chip memory.
@@ II-E A Motivating Case Study: Attention + Linear Projection
The pattern of data size variation is first increase then decrease. Dataflow execution instantiates PEs for each task. There are three ways to stream and cache data across tasks, illustrated in Figure 1: outer product with aligned and unaligned production rates of $Q$ and $K$ (Figure 1(a) and (b)), and inner product with unaligned production rates (Figure 1(c)). Inner products with aligned production rates require caching of both $Q$ and $K$, which is infeasible. Each case incurs compute resource idling due to either varied workload size of $A$, back-pressure from streaming, or early completion of tasks.
@@ III Inter-Task Auto-Reconfigurable Accelerator
Unlike other reconfigurable [20, 18, 30, 17, 21] and overlay accelerators [31, 32], which prioritize abstraction and fast compilation, InTAR focuses on reconfiguration for optimization. The reconfiguration schedule in InTAR is determined at circuit design time rather than relying on post-design external commands. For every application, a static schedule is created based on the specific application and resource constraints, and the circuit is then tailored accordingly. This approach reduces reliance on general-purpose control units and cross-bars, retaining only essential reconfiguration logic and interconnects. The result is low resource utilization for reconfiguration controls (49% LUT reduction for a design with 3×3$\times$3 × throughput compared to [20] for multiple GEMMs), high clock frequency (2.41×2.41$\times$2.41 × that of [20]), and low reconfiguration latency ($10^{6}\times$ faster than DPR [21], and comparable to [20]) compared with FPGA-based CGRA and DPR approaches. Additionally, InTAR handles reconfiguration at the task level, enabling rapid and convenient development with high-level tools like HLS. This makes InTAR well-suited for FPGAs to adapt quickly to new DNN applications. Lastly, InTAR’s reconfiguration scope is not limited to specific computations. Table II compares InTAR with prior DNN accelerator works.
@@ III-A Execution Modes of InTAR
Figure 2: Left: Architecture template of InTAR. Compute cores (CC) compute linear operations (e.g., GEMM, ConvNet), and SFUs compute non-linear operations (e.g., softmax, GeLU). Each CC contains a scratchpad memory, a reconfigurable MAC unit array, a reduction unit, and a data movement control unit. Dashed lines indicate the candidates for connection between CCs and SFUs. Right: example architectures for each execution mode within the template, with $n,m=2$. For a schedule, we will compose the architectures of the required modes to keep only the necessary interconnects and logic.
@@
  • Task-pipeline: Same as dataflow execution with pipelined dependent tasks. Computational resources are partitioned proportionally to the workloads. Data are streamed from one task to another using FIFOs.
  • Task-parallel: Same as dataflow execution with independent tasks running in parallel. The resource efficiency is the same as the sequential mode, but it may have a better locality for subsequent tasks (e.g., for the example in upper left of Figure 1, if calculating $V^{\prime}$ and $O$ are pipelined, then running $Q$ and $V$ in parallel and assigning $V$ to the PE that will compute $V^{\prime}$ avoids data aggregation overhead from other PEs).
Figure 1(f) illustrates the task assignment and timeline of InTAR for the case study in Section II-E. PEs 1 and 2 handle both attention computation and linear projections. Depending on the subsequent operations, InTAR can choose either task-parallel or sequential mode for $Q$ and $V$. If the scheduler decides to serialize computing $V^{\prime}$ and $O$, then the sequential mode is more suitable to allow PEs 1 and 2 to execute cooperatively. If the scheduler pipelines computations of $V^{\prime}$ and $O$, then task-parallel mode is preferred to improve locality. Both choices will not affect the latency but will determine the data movement overhead of the following tasks. After calculating $Q$ and $V$, InTAR switches to task-pipeline mode to compute $K$ and $A$, since storing $Q$ consumes all scratchpad memory.
@@ III-B Architecture Template
  • Reconfigurable PE Array (PEA): a multi-dimensional array of PEs that allows changing inputs and precisions. Data communication between PEs is either in a systolic array style or by broadcasting the data to all PEs. Each PE produces the complete or partial result of the linear operation.
@@
When switching execution modes, a global instruction reader will read both the configuration instructions from a static buffer and the input size passed as the kernel argument. It will then modify part of each instruction related to the input size (e.g., the loop bounds) and send it to the CCs. The reconfigurable PEA and data movement control will read the instructions and propagate them among the CCs. The instruction sequence is determined by the static schedule of mode switching at the circuit design time. Each instruction contains a stage index and several loop bounds (e.g., reconfiguring the three loop bounds for GEMMs). A stage refers to the execution of a single or a group of tasks that follows one of the execution modes, seen in Figure 1(f), e.g., computing $Q$ and $V$ in stage 1 and $K$ and $A$ in stage 2. Each CC has a specified behavior for each stage, orchestrated by multiplexers with stage indices as the select signals. Notice that the stage-specific behaviors vary from one application to another. Thus, InTAR only needs to handle the reconfigurations defined in that application-specific schedule. The reconfiguration overhead consists of four parts: reading, modifying, sending, and decoding instructions to generate signals for multiplexers in the CCs. Each of these is done in a single cycle. Thus, reconfigurations have a 4-cycle overhead (10 to 20 ns, depending on frequency), which is negligible compared to modern DNN inference latency.
@@ IV Designing Accelerators in InTAR
The first step in designing InTAR accelerators is performing design space exploration (DSE). Given a dataflow graph (DFG) and device constraints, one searches through various combinations of architecture parameters and task execution modes (see Section III) to minimize total latency.
@@ IV-A Finding Schedule and Architecture Template Parameters
In this work, we first topologically sort the tasks based on dependencies in the DFG and apply a heuristic-based DSE:
@@
Then, we analyze the generated schedule and the hardware configurations and determine architecture template parameters with the following heuristics: The schedule infers the instantiation of the SFUs and FIFO connections. The column count of CCs equals the maximum number of tasks of all stages, and the row count is the number of dies on the FPGA necessary to improve floorplanning. We compute the PEA dimensions to maximize utilization with limited cross-die communications and calculate the memory ports needed. Finally, the scratchpad memory is the maximum memory size required over all stages. For applications with variable input size, we schedule based on the maximum possible size and employ batching [33] to prevent inefficiency of the design for small data size. Although the heuristics are not guaranteed optimal, we show that our schedule results improved performance consistently (see Sections V, VI). We leave the development of an optimized DSE engine as future work.